# The resummation model in FLAMINGO: precisely predicting matter power suppression from observed halo baryon fractions

### v1

@@ 2.2 Resummation model
The prototypical resummation model for predicting the suppression of the matter power spectrum by baryonic feedback was presented in van Loon & van Daalen (2024). In the current work, we extend and improve the model, and apply it to FLAMINGO. For completeness, we describe the full model here.
@@ 2.2.1 Accounting for mass loss
Summarising what we have presented so far, this model takes dark-matter-only halo mass fractions, linear biases and (cross-)power spectra, and combines them with the mean fraction of mass retained by haloes that have undergone galaxy formation, relative to their dark-matter-only equivalent, to predict the change in the total matter power spectrum. This mean fraction of mass retained, $f_{\mathrm{ret},i,\Delta}$, can be calculated as the total mass ratio of (matched) haloes in a hydrodynamical simulation and its DMO equivalent. Observationally, $f_{\mathrm{ret},i,\Delta}$ can in principle be derived from the mean observed baryon fractions of haloes of a certain mass, $\bar{f}_{\mathrm{b},i,\Delta}$, relative to the cosmic baryon fraction $\Omega_{\mathrm{b}}/\Omega_{\mathrm{m}}$. Specifically, under the assumption that all matter removed from the halo was baryonic matter, one would find the retained total matter fraction to be:
@@
However, in reality the dark matter will respond gravitationally to the loss of baryonic mass, and additional mass will be lost as the halo relaxes (see also §4.1). van Loon & van Daalen (2024) noted that a linear relation between the retained mass fraction and the halo baryon fraction, with a slope above unity, reproduced the results of the cosmo-OWLS (Le Brun et al., 2014) and BAHAMAS (McCarthy et al., 2017) simulations well. However, the larger volumes of FLAMINGO allow us to explore this relation with far better statistics, and to larger mass scales. We find that the following functional form is able to accurately fit all FLAMINGO haloes with masses $M_{\mathrm{h},i,\mathrm{500c}}\gtrsim 10^{12}\,\mathrm{M}_{\odot}$, for different values of $\Delta$:
@@ 2.2.2 Combining overdensity regions
Say that we measure the total (retained) mass and baryon fraction of a sample of haloes for both $\Delta=\mathrm{500c}$ and $\Delta=\mathrm{200m}$. We can then calculate the retained fraction for each region with equation (9). The net fraction of the “original” DMO mass that has been removed from $R_{\mathrm{500c}}$, but not from $R_{\mathrm{200m}}$, is then $f_{\mathrm{ret},i,\mathrm{200m}}-f_{\mathrm{ret},i,\mathrm{500c}}$. As shown in van Loon & van Daalen (2024), this fraction is quite significant for massive haloes. We will refer to the region $R_{\mathrm{500c}}\,<r<\,R_{\mathrm{200m}}$ as the halo annulus, and will indicate it with subscript AA. [^10] Note that $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$ can be replaced by any other two regions in which the halo baryon fraction was measured, such as $R_{\mathrm{200c}}$ and $R_{\mathrm{50c}}$.
One thing that was not taken into account in van Loon & van Daalen (2024), is that conservation of the product of mass and bias does not just apply to mass that is added to the non-halo matter, but generally, to mass moved between any two regions. While we have checked that for the regions we have chosen here the distinction between mass-bias conservation or simply mass conservation leads to negligible differences, as the $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$ regions have a virtually identical bias, we still adapt equations (10) and (11) from van Loon & van Daalen (2024) to include the bias correctly. The cross-power contribution of matter in these regions is then given by:
@@ 2.2.3 Pushing to smaller scales
On small scales, the matter power spectrum increases more steeply in baryonic simulations than in dark-matter-only simulations, due to gas cooling into stars, forming galaxies, and the contraction of the inner dark matter halo in response. We can crudely include the effect of this on larger scales ($k\,{\sim}\,{10}\,h\mathrm{\,Mpc^{-1}}$) in the resummation model as follows. First, we determine the cross power spectrum of halo centres with all matter, $P_{\mathrm{hc},i}$, in the DMO simulation, where each halo centre is represented by a single particle with mass equal to the $M_{\mathrm{500c}}$ mass of the halo. We then make the approximation that all stellar mass within a region, e.g. $R_{\mathrm{2500c}}$, is located at its centre. Under the (implicit) additional assumption that the stellar mass of a halo is roughly proportional to its $M_{\mathrm{500c}}$ within a mass bin, we can then rescale the halo centre power spectrum to the stellar mass contribution to the power spectrum:
@@ 2.2.4 Model summary
A Python package called resummation implementing these steps, along with the necessary DMO cross spectra and other quantities for the FLAMINGO simulations, will be made available upon acceptance of this manuscript for publication. The code calculates power suppression predictions in $1\,\mathrm{ms}$ for a given cosmology, allowing for fast iterations to marginalise over input halo baryon fractions.
@@ 3.1 Retained mass vs baryon fraction
By matching haloes between each hydrodynamical FLAMINGO simulation and its DMO equivalent (following the procedure described in §3.2 of Elbers et al., 2025), we can calculate the mean retained mass fraction in each halo mass bin as the ratio of the overdensity mass in the baryonic simulation and in DMO. The mean baryon fractions in the hydrodynamical haloes, $f_{\mathrm{bc},i,\Delta}$, can then be mapped to these retained mass fractions $f_{\mathrm{ret},i,\Delta}$. In practice, the baryon fractions are measured at a total observed halo mass, but to apply the resummation model we need to know the retained mass fraction at a given DMO mass. This is not an issue, since the retained mass fraction itself gives us exactly the ratio between these masses. However, to eliminate this extra step, we derive the mappings between the observed baryon fraction and both the retained mass fraction for those same haloes (i.e. at the DMO mass that those haloes would have had in a DMO universe) and the retained mass fraction for haloes that have a total mass in DMO equal to the observed mass (i.e. mapping baryon fractions of haloes with mass $M_{\mathrm{obs},i}$ to retained mass fractions of haloes with mass $M_{\mathrm{DMO},j}=M_{\mathrm{obs},i}$). We use the latter relation for all later results. Since for this version of the relation, the overdensity mass of any observed halo and the DMO halo it is mapped to are the same, their overdensity radii are equal as well, and thus we do not need to assume that the ejection of matter leaves the overdensity radius unchanged in our application of the model.
The results of these mappings are shown in Fig. 2. Note that we present these mappings in terms of the corrected baryon fractions, $f_{\mathrm{bc}}$ (see equation 8), which takes out some dependence on the cosmic baryon fraction. The corrected baryon fraction is defined such that it is equal to the retained mass fraction when only baryons are removed from haloes, relative to DMO – that is, if there is no dark matter response or halo relaxation. This one-to-one relation is shown as a dotted line. Every point represents one $0.5\,\mathrm{dex}$ $M_{\mathrm{500c}}$ halo mass bin, with minimum mass $M_{\mathrm{500c,DMO}}=10^{12}\,\mathrm{M_{\odot}}$, up to the highest-mass haloes present in the simulation volume. Every colour represents a different hydrodynamical simulation, following the colour scheme and naming convention of Schaye et al. (2023).
@@
Strikingly, the correlation between the retained mass fraction and the corrected baryon fraction shown in Fig. 2 is very tight. This shows that the retained mass fraction can be predicted precisely from observed baryon fractions, independent of cosmology or feedback strength. This is true regardless of whether we map to the retained mass fractions of the same haloes (left panel) or those of haloes with the same total mass (right panel). The deviations from the one-to-one relation are also clear: as the (corrected) baryon fraction decreases, the retained mass fraction drops further below the dotted line. This is expected: the more mass is removed, the stronger the dark matter response. Additionally, by removing baryons the overdensity radius shrinks, lowering the overdensity mass even without ejecting matter (e.g. Velliscig et al., 2014, see also §4.1). At high baryon fractions, however, the retained mass fractions exceed the dotted line. This implies that for these halo masses, halo contraction due to gas cooling and star formation is more important than mass ejection.
@@
The highest-mass haloes of each simulation are found in the top right of each panel, where feedback has not been strong enough – relative to the halo’s potential – to eject much gas beyond $R_{\mathrm{200m}}$, or at least not significantly more gas than was gained by inflow (see e.g. Mitchell & Schaye, 2022). As the halo mass decreases, haloes move down the relation to the left. The speed at which they move down the relation depends on the simulation parameters. However, interestingly, when the halo mass becomes low enough for AGN feedback to become ineffective, and supernova feedback begins to dominate ($M_{\mathrm{500c}}\lesssim 10^{13}\,\mathrm{M_{\odot}}$), the points move back up the same relation, rather than deviating from it. This shows that this mapping is not specific to AGN feedback, but applies more broadly. We note that we did not cut off the x-axis at an arbitrary value: no halo mass bin in any of the simulations showed $f_{\mathrm{bc},i,\Delta}<0.88$, for any $\Delta$. The scatter between simulations gradually increases with decreasing halo mass, which may be caused by the numerical resolution as low-mass haloes are more poorly resolved (as mentioned previously, the lowest-mass haloes we include here have only $\sim 100$ DM particles).
@@ 3.2 Cross-power model
In both panels of Fig. 1, we see the same trend with mass: low-mass haloes contribute very little to the total signal, but the contribution grows with mass as both the halo bias and the total mass inside the overdensity do – until the contribution peaks for $M_{\mathrm{500c}}\sim 10^{14}\,\mathrm{M_{\odot}}$, after which the bias continues to increase but the mass fraction strongly declines. This is in line with the findings of van Loon & van Daalen (2024) for cosmo-OWLS and BAHAMAS, though we are able to extend this to larger mass scales and with less noise, due to the much larger volume of FLAMINGO.
@@ 3.3 Suppression signals from the resummation model
Figure 5: Model predictions compared to the true suppression signal for FLAMINGO simulation L1_m9, which is shown as a solid red line. Red dotted lines show the $\pm 1$ percentage-point band around the true suppression signal. The van Daalen et al. (2020, VD20) relation, based on the mean baryon fraction in haloes of mass $M_{\mathrm{500c}}\approx 10^{14}\,\mathrm{M_{\odot}}/h$ and fit to simulation results up to $k\,{=}\,{1}\,h\mathrm{\,Mpc^{-1}}$, is shown as a dashed cyan line for comparison. The results from the resummation model are shown as dot-dashed lines. Using only the mean halo baryon fractions within $R_{\mathrm{200m}}$ as input reproduces the true suppression signal within $1$ percentage point up to $k\,{\approx}\,{3}\,h\mathrm{\,Mpc^{-1}}$, while using only those within $R_{\mathrm{500c}}$ follows the true signal down to smaller scale but underpredicts the amount of suppression on large scales. However, combining baryon fractions from both regions, as described in §2.2.2, allows us to reproduce the true signal to within $1$ percentage point up to $k\,{\approx}\,{10}\,h\mathrm{\,Mpc^{-1}}$.
Figure 6: The results of applying the resummation model to all baryonic FLAMINGO variations, using mean halo baryon fractions measured from these simulations within both $R_{\mathrm{200m}}$ and $R_{\mathrm{500c}}$ for haloes $M_{\mathrm{500c}}\geq 10^{12.5}\,\mathrm{M_{\odot}}$. The top panels show the suppression signals, while the bottom panels show the ratio of the model predictions to the true suppression signals of the simulations. Variations with the default D3A cosmology are shown in the left panels, those with other cosmologies in the right (with L1_m9 and L2p8_m9 added as a reference, which are nigh indistinguishable from each other). The true suppression signals of the FLAMINGO simulations (solid lines) are all accurately reproduced by the resummation model (dashed lines) down to the scale on which the suppression signal starts to turn up, caused by the formation of galaxies in the centre of haloes and the contraction of the inner halo in response, which is not captured by the default model. Strikingly, the model even accurately reproduces the relative power spectrum in a simulation without any cooling or feedback (black line). We emphasise that the model has zero free parameters.
In Fig. 5, we compare the true suppression signal for the L1_m9 simulation, $P_{\mathrm{hydro}}(k)/P_{\mathrm{DMO}}(k)$, to different predictions. These suppression signals have been smoothed by convolving with a Gaussian with a standard deviation equal to two data points for clarity. We show the true suppression signal as a solid red line, with dotted lines showing a $\pm 1$ percentage-point band around it – ideally, our predictions are within this band, at least up to $k\,{=}\,{10}\,h\mathrm{\,Mpc^{-1}}$. As a comparison, a dashed cyan line shows the prediction of van Daalen et al. (2020, VD20), which is based only on the mean baryon fraction of $M_{\mathrm{500c}}\sim 10^{14}\,\mathrm{M_{\odot}}$ haloes. While this accurately predicts the suppression signal on scales where this halo contribution is the dominant one, it fails for scales $k\,{\gtrsim}\,{2}\,h\mathrm{\,Mpc^{-1}}$, for reasons explained in Section 1.
@@
In fact, for the FLAMINGO simulation without gas cooling, which has neither star formation nor feedback but does show the effects of gas pressure relative to DMO, the model reproduces the true signal to within $1$ percentage point on all scales shown, $k\,{\leq}\,{25}\,h\mathrm{\,Mpc^{-1}}$ (black lines). This is striking, as the shape of its suppression signal is completely different from that of simulations with strong feedback, and therefore not easily captured by existing fitting functions (e.g. Schaller et al., 2025). The resummation model, which does not rely on any particular functional form as a function of $k$, but instead rescales DMO cross spectra, is much more flexible.

### v2

@@ 2.2 Resummation model
The prototypical resummation model for predicting the suppression of the matter power spectrum by baryonic feedback was presented in vanLoonvanDaalen2024. In the current work, we extend and improve the model, and apply it to FLAMINGO. For completeness, we describe the full model here, and note where we deviate from the model description of vanLoonvanDaalen2024, which is otherwise reproduced below. On top of this, compared to this previous work we have added some clarifications and summaries.
@@ 2.2.1 Accounting for mass loss
Summarising what we have presented so far, this model takes dark-matter-only halo mass fractions, linear biases and (cross-)power spectra, and combines them with the mean fraction of mass retained by haloes that have undergone galaxy formation, relative to their dark-matter-only equivalent, to predict the change in the total matter power spectrum. This mean fraction of mass retained, $f_{\mathrm{ret},i,\Delta}$, can be calculated as the total mass ratio of (matched) haloes in a hydrodynamical simulation and its DMO equivalent. [^8] The halo overdensity radii of hydrodynamical and DMO haloes will in general be different – however, we will demonstrate in §3.1 that this can be fully accounted for. Observationally, $f_{\mathrm{ret},i,\Delta}$ can in principle be derived from the mean observed baryon fractions of haloes of a certain mass, $\bar{f}_{\mathrm{b},i,\Delta}$, relative to the cosmic baryon fraction $\Omega_{\mathrm{b}}/\Omega_{\mathrm{m}}$. Specifically, under the assumption that all matter removed from the halo was baryonic matter, one would find the retained total matter fraction to be:
@@
However, in reality the dark matter will respond gravitationally to the loss of baryonic mass, and additional mass will be lost as the halo relaxes (see also §4.1). vanLoonvanDaalen2024 noted that a linear relation between the retained mass fraction and the halo baryon fraction, with a slope above unity, reproduced the results of the cosmo-OWLS (LeBrun2014) and BAHAMAS (McCarthy2017) simulations well. However, the larger volumes of FLAMINGO allow us to explore this relation with far better statistics, and to larger mass scales. We find that the following functional form is able to accurately fit all FLAMINGO haloes with masses $M_{\mathrm{h},i,\mathrm{500c}}\gtrsim 10^{12}\,\mathrm{M}_{\odot}$, for different values of $\Delta$:
@@ 2.2.2 Combining overdensity regions
Say that we measure the total (retained) mass and baryon fraction of a sample of haloes for both $\Delta=\mathrm{500c}$ and $\Delta=\mathrm{200m}$. We can then calculate the retained fraction for each region with equation (9). The net fraction of the “original” DMO mass that has been removed from $R_{\mathrm{500c}}$, but not from $R_{\mathrm{200m}}$, is then $f_{\mathrm{ret},i,\mathrm{200m}}-f_{\mathrm{ret},i,\mathrm{500c}}$. As shown in vanLoonvanDaalen2024, this fraction is quite significant for massive haloes. We will refer to the region $R_{\mathrm{500c}}\,<r<\,R_{\mathrm{200m}}$ as the halo annulus, and will indicate it with subscript AA. [^11] Note that $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$ can be replaced by any other two regions in which the halo baryon fraction was measured, such as $R_{\mathrm{200c}}$ and $R_{\mathrm{50c}}$.
One thing that was not taken into account in vanLoonvanDaalen2024, is that conservation of the product of mass and bias does not just apply to mass that is added to the non-halo matter, but generally, to mass moved between any two regions. While we have checked that for the regions we have chosen here the distinction between mass-bias conservation or simply mass conservation leads to negligible differences, as the $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$ regions have a virtually identical bias, we still adapt equations (10) and (11) from vanLoonvanDaalen2024 to include the bias correctly. The cross-power contribution of matter in these regions is then given by:
@@ 2.2.3 Pushing to smaller scales
On small scales, the matter power spectrum increases more steeply in baryonic simulations than in dark-matter-only simulations, due to gas cooling into stars, forming galaxies, and the contraction of the inner dark matter halo in response. Going beyond what was done in vanLoonvanDaalen2024, we can crudely include the effect of this on larger scales ($k\,{\sim}\,{10}\,h\mathrm{\,Mpc^{-1}}$) in the resummation model as follows. First, we determine the cross power spectrum of halo centres with all matter, $P_{\mathrm{hc},i}$, in the DMO simulation, where each halo centre is represented by a single particle with mass equal to the $M_{\mathrm{500c}}$ mass of the halo. We then make the approximation that all stellar mass within a region, e.g. $R_{\mathrm{2500c}}$, is located at its centre. Under the (implicit) additional assumption that the stellar mass of a halo is roughly proportional to its $M_{\mathrm{500c}}$ within a mass bin, we can then rescale the halo centre power spectrum to the stellar mass contribution to the power spectrum:
@@ 2.2.4 Model summary
A Python package called resummation implementing these steps, along with the necessary DMO cross spectra and other quantities for the FLAMINGO simulations, is publicly available (see Data Availability). The code calculates power suppression predictions in $1\,\mathrm{ms}$ for a given cosmology, allowing for fast iterations to marginalise over input halo baryon fractions.
@@ 3.1 Retained mass vs baryon fraction
By matching haloes between each hydrodynamical FLAMINGO simulation and its DMO equivalent (following the procedure described in §3.2 of Elbers2025), we can calculate the mean retained mass fraction in each halo mass bin as the ratio of the overdensity mass in the baryonic simulation and in DMO. The mean baryon fractions in the hydrodynamical haloes, $f_{\mathrm{bc},i,\Delta}$, can then be mapped to these retained mass fractions $f_{\mathrm{ret},i,\Delta}$. In practice, the baryon fractions are measured at a total observed halo mass, but to apply the resummation model we need to know the retained mass fraction at a given DMO mass. This is not an issue, since the retained mass fraction itself gives us exactly the ratio between these masses. However, to eliminate this extra step, we derive the mappings between the observed baryon fraction and both the retained mass fraction for those same haloes (i.e. at the DMO mass that those haloes would have had in a DMO universe) and the retained mass fraction for haloes that have a total mass in DMO equal to the observed mass (i.e. mapping baryon fractions of haloes with mass $M_{\mathrm{obs},i}$ to retained mass fractions of haloes with mass $M_{\mathrm{DMO},j}=M_{\mathrm{obs},i}$). We use the latter relation for all later results. Since for this version of the relation, the overdensity mass of any observed halo and the DMO halo it is mapped to are the same, their overdensity radii are equal as well, and thus we do not need to assume that the ejection of matter leaves the overdensity radius unchanged in our application of the model.
The results of these mappings are shown in Fig. 2. Note that we present these mappings in terms of the corrected baryon fractions, $f_{\mathrm{bc}}$ (see equation 8), which takes out some dependence on the cosmic baryon fraction. The corrected baryon fraction is defined such that it is equal to the retained mass fraction when only baryons are removed from haloes, relative to DMO – that is, if there is no dark matter response or halo relaxation. This one-to-one relation is shown as a dotted line. Every point represents one $0.5\,\mathrm{dex}$ $M_{\mathrm{500c}}$ halo mass bin, with minimum mass $M_{\mathrm{500c,DMO}}=10^{12}\,\mathrm{M_{\odot}}$, up to the highest-mass haloes present in the simulation volume. Every colour represents a different hydrodynamical simulation, following the colour scheme and naming convention of Schaye2023.
@@
Strikingly, the correlation between the retained mass fraction and the corrected baryon fraction shown in Fig. 2 is very tight. This shows that the retained mass fraction can be predicted accurately from observed baryon fractions, independent of cosmology or feedback strength. This is true regardless of whether we map to the retained mass fractions of the same haloes (left panel) or those of haloes with the same total mass (right panel). The deviations from the one-to-one relation are also clear: as the (corrected) baryon fraction decreases, the retained mass fraction drops further below the dotted line. This is expected: the more mass is removed, the stronger the dark matter response. Additionally, by removing baryons the overdensity radius shrinks, lowering the overdensity mass even without ejecting matter (e.g. Velliscig2014, see also §4.1). At high baryon fractions, however, the retained mass fractions exceed the dotted line. This implies that for these halo masses, halo contraction due to gas cooling and star formation is more important than mass ejection.
@@
The highest-mass haloes of each simulation are found in the top right of each panel, where feedback has not been strong enough – relative to the halo’s potential – to eject much gas beyond $R_{\mathrm{200m}}$, or at least not significantly more gas than was gained by inflow (see e.g. MitchellSchaye2022). As the halo mass decreases, haloes move down the relation to the left. The speed at which they move down the relation depends on the simulation parameters. However, interestingly, when the halo mass becomes low enough for AGN feedback to become ineffective, and supernova feedback begins to dominate ($M_{\mathrm{500c}}\lesssim 10^{13}\,\mathrm{M_{\odot}}$), the points move back up the same relation, rather than deviating from it. This shows that this mapping is not specific to AGN feedback, but applies more broadly. We note that we did not cut off the x-axis at an arbitrary value: no halo mass bin in any of the simulations showed $f_{\mathrm{bc},i,\Delta}<0.88$, for any $\Delta$. The scatter between simulations gradually increases with decreasing halo mass, which may be caused by the numerical resolution as low-mass haloes are more poorly resolved (as mentioned previously, the lowest-mass haloes we include here have only $\sim 100$ DM particles).
@@ 3.2 Cross-power model
In both panels of Fig. 1, we see the same trend with mass: low-mass haloes contribute very little to the total signal, but the contribution grows with mass as both the halo bias and the total mass inside the overdensity do – until the contribution peaks for $M_{\mathrm{500c}}\sim 10^{14}\,\mathrm{M_{\odot}}$, after which the bias continues to increase but the mass fraction strongly declines. This is in line with the findings of vanLoonvanDaalen2024 for cosmo-OWLS and BAHAMAS, though we are able to extend this to larger mass scales and with less noise, due to the much larger volume of FLAMINGO.
@@
Figure 5: Model predictions compared to the true suppression signal for FLAMINGO simulation L1_m9, which is shown as a solid red line. Red dotted lines show the $\pm 1$ percentage-point band around the true suppression signal. The vanDaalen2020 relation, based on the mean baryon fraction in haloes of mass $M_{\mathrm{500c}}\approx 10^{14}\,\mathrm{M_{\odot}}/h$ and fit to simulation results up to $k\,{=}\,{1}\,h\mathrm{\,Mpc^{-1}}$, is shown as a dashed cyan line for comparison. The results from the resummation model are shown as dot-dashed lines. Using only the mean halo baryon fractions within $R_{\mathrm{200m}}$ as input reproduces the true suppression signal within $1$ percentage point up to $k\,{\approx}\,{3}\,h\mathrm{\,Mpc^{-1}}$, while using only those within $R_{\mathrm{500c}}$ follows the true signal down to smaller scale but underpredicts the amount of suppression on large scales. However, combining baryon fractions from both regions, as described in §2.2.2, allows us to reproduce the true signal to within $1$ percentage point up to $k\,{\approx}\,{10}\,h\mathrm{\,Mpc^{-1}}$.
Figure 6: The results of applying the resummation model to all baryonic FLAMINGO variations, using mean halo baryon fractions measured from these simulations within both $R_{\mathrm{200m}}$ and $R_{\mathrm{500c}}$ for haloes $M_{\mathrm{500c}}\geq 10^{12.5}\,\mathrm{M_{\odot}}$. The top panels show the suppression signals, while the bottom panels show the ratio of the model predictions to the true suppression signals of the simulations. Variations with the default D3A cosmology are shown in the left panels, those with other cosmologies in the right (with L1_m9 and L2p8_m9 added as a reference, which are nigh indistinguishable from each other). The true suppression signals of the FLAMINGO simulations (solid lines) are all accurately reproduced by the resummation model (dashed lines) down to the scale on which the suppression signal starts to turn up, caused by the formation of galaxies in the centre of haloes and the contraction of the inner halo in response, which is not captured by the default model. Strikingly, the model even accurately reproduces the relative power spectrum in a simulation without any cooling or feedback (black line). We emphasise that the model uses zero free parameters to transform given halo baryon fractions into a predicted suppression signal.
@@ 3.3 Suppression signals from the resummation model
In Fig. 5, we compare the true suppression signal for the L1_m9 simulation, $P_{\mathrm{hydro}}(k)/P_{\mathrm{DMO}}(k)$, to different predictions. These suppression signals have been smoothed by convolving with a Gaussian with a standard deviation equal to two data points for clarity. We show the true suppression signal as a solid red line, with dotted lines showing a $\pm 1$ percentage-point band around it – ideally, our predictions are within this band, at least up to $k\,{=}\,{10}\,h\mathrm{\,Mpc^{-1}}$. As a comparison, a dashed cyan line shows the prediction of vanDaalen2020, which is based only on the mean baryon fraction of $M_{\mathrm{500c}}\sim 10^{14}\,\mathrm{M_{\odot}}$ haloes. While this accurately predicts the suppression signal on scales where this halo contribution is the dominant one, it fails for scales $k\,{\gtrsim}\,{2}\,h\mathrm{\,Mpc^{-1}}$, for reasons explained in Section 1.
@@
In fact, for the FLAMINGO simulation without gas cooling, which has neither star formation nor feedback but does show the effects of gas pressure relative to DMO, the model reproduces the true signal to within $1$ percentage point on all scales shown, $k\,{\leq}\,{25}\,h\mathrm{\,Mpc^{-1}}$ (black lines). This is striking, as the shape of its suppression signal is completely different from that of simulations with strong feedback, and therefore not easily captured by existing fitting functions (e.g. Schaller2025). The resummation model, which does not rely on any particular functional form as a function of $k$, but instead rescales DMO cross spectra, is much more flexible.